OpenStudy (anonymous):

Base and height of triangle whose height is four more than seven times its base. Whats the base and height? Thanks :)

OpenStudy (anonymous):

h = 7a + 4 a = (h-4)/7

OpenStudy (anonymous):

whats the seven?

OpenStudy (diyadiya):

height is four more than 7 times its base so h= 7b+4 where b is base so 7b=h-4 b=(h-4)/7

OpenStudy (anonymous):

/7 is what i meant :) but thank you for that information

OpenStudy (diyadiya):

7 is given in the question :) "seven times its base " :)

OpenStudy (anonymous):

How do i get the answers, do i solve both?

OpenStudy (diyadiya):

i hope this is the final answer

OpenStudy (anonymous):

which one is the final answer :) is the answer just how its set up?

OpenStudy (diyadiya):

\[h= 7b+4 \]\[b= \frac{(h-4)}{7}\]

OpenStudy (anonymous):

Thank you :) Can you help me with two more? Then I'll be done for now. I'm just on a less time because I got up late and I have to work on another project soon. But first I have to get this math out of the way, my teacher makes us have perfectly correct answers

OpenStudy (diyadiya):

I'll try

OpenStudy (diyadiya):

wait 5min

OpenStudy (anonymous):

ok :D thanks so much, ill be right back...

OpenStudy (anonymous):

one leg of a right triangle is 6 millimeters shorter than longer leg, and hypotenuse is 6 millimeters longer than longer leg. Find lengths of sides of triangle.

OpenStudy (anonymous):

so i need longer, shorter, and hypo :)

OpenStudy (diyadiya):

Longleg=x shorter leg= x-6 Hypo=x+6 (x+6)^2=(x-6)^2+x^2 x^2+12x+36=x^2-12x+36+x^2 x^2-24x=0 x(x-24)=0 x=0 or x=24

OpenStudy (anonymous):

Yay thanks so much! One more question and I'm done...be right back :D

OpenStudy (diyadiya):

ok :)

OpenStudy (anonymous):

oh which is longer etc?

OpenStudy (anonymous):

It wants three answers

OpenStudy (diyadiya):

i've , mentioned everything can you tell me which is longer , shorter & hypotenuse?

OpenStudy (anonymous):

hypo: 30 longer:24 short:19?

OpenStudy (diyadiya):

is short one 19?

OpenStudy (diyadiya):

24-6=18 :)

OpenStudy (anonymous):

18? ok thanks :D Ill be back in a sec

OpenStudy (anonymous):

Ok same problem but 1 mill instead of six :( My computer kicked me off and made me a new problem, I'm sorry

OpenStudy (diyadiya):

i didnt get you,what happend?

OpenStudy (anonymous):

The problem reset

OpenStudy (anonymous):

can you solve it with 1 instead of six used? mill wise :)

OpenStudy (diyadiya):

can you try out i'll help you

OpenStudy (anonymous):

ok just a sec :)

OpenStudy (anonymous):

I tried :( I just messed the whole thing up, the dumb x^2 confuses me

OpenStudy (diyadiya):

(x+1)^2=(x-1)^2+x^2 x^2+2x+1=x^2-2x+1+x^2 x^2-4x=0 x(x-4)=0 x=0 or x=4 longest=4 shortest =4-1=3 hypo=4+1=5

OpenStudy (anonymous):

oh Diya thanks so much :)

OpenStudy (diyadiya):

Welcome =)

OpenStudy (anonymous):

I have one more now ^_^ <3